OpenStudy (anonymous):

If r(t)=cos(7t)i+sin(7t)j+2tk, compute the tangential and normal components of the acceleration vector.

OpenStudy (amistre64):

i know a paul h dunn ...

OpenStudy (amistre64):

just did one of these, or tried to :)

OpenStudy (amistre64):

\[a_t=\frac{r'*r''}{|r'|}\] \[a_n=\frac{|r'x\ r''|}{|r'|}\]

OpenStudy (anonymous):

I have the 1st part done, I know its 0 I cant figure out the second

OpenStudy (amistre64):

http://tutorial.math.lamar.edu/Classes/CalcIII/Velocity_Acceleration.aspx is what I got to help refresh me memory

OpenStudy (amistre64):

r(t)=<cos(7t),sin(7t),2t> r'(t)=<-7sin(7t),7cos(7t),2> r''(t)=<-49cos(7t),-49sin(7t),0>

OpenStudy (amistre64):

< -7sin(7t), 7cos(7t),2> <-49cos(7t),-49sin(7t),0> ------------------------- i = 98sin(7t) j = -98cos(7t) k = 7^3 sin^2(t) + 7^3 cos^2(t) = 7^3 r'xr'' = <98sin(7t),-98cos(7t), 7^3> |r'xr''| = sqrt( 98^2sin^2(7t)+98^2cos^2(7t)+7^6) = sqrt( 98^2+7^6)

OpenStudy (anonymous):

Thats correct, I have gotten the same thing so far

OpenStudy (anonymous):

I jut have to put it all together haha

OpenStudy (amistre64):

as far as i can tell; you divide that by the magnitude of r'

OpenStudy (amistre64):

r' = <-7sin(7t),7cos(7t),2> |r'| = sqrt(49sin^2(7t)+49cos^2(7t)+4) = sqrt(49+4) = sqrt(53)

OpenStudy (amistre64):

i get 49 ...

OpenStudy (anonymous):

that what I am thinking, is the mag of r prime not ((7^2)+(7^2++(2^2))

OpenStudy (anonymous):

Wow im dumb haha

OpenStudy (anonymous):

So you in college haha?

OpenStudy (amistre64):

not that I can algebra out: \[\sqrt{49sin^2(7t)+49cos^2(7t)+4}\] \[\sqrt{49(sin^2(7t)+cos^2(7t))+4}\] \[\sqrt{49(1)+4}\] \[\sqrt{49+4}\]

OpenStudy (amistre64):

yep, in college

OpenStudy (anonymous):

thats correct! I typed in the the computer homework set

OpenStudy (anonymous):

I really apprectiate your help

OpenStudy (amistre64):

youre welcome, and good luck :)

OpenStudy (amistre64):

someplace in florida; home of the lions or something like that